OpenStudy (anonymous):

in a 45 45 90 triangle if the legs each have a length of 4 then what is the length of the hypotenuse

OpenStudy (anonymous):

\[4^2+4^2=c^2\] that becomes 16+16=c^2 32=c^2 square root both sides and you get \[\sqrt{32}\] if yu want it simplified its \[4\sqrt{2}\]

OpenStudy (anonymous):

thanks!